# CADRE: Customizable Assurance of Data Readiness in Privacy-Preserving Federated Learning

### v1

@@ Abstract
Privacy-Preserving Federated Learning (PPFL) is a decentralized machine learning approach where multiple clients train a model collaboratively. PPFL preserves privacy and security of the client’s data by not exchanging it. However, ensuring that data at each client is of high quality and ready for federated learning (FL) is a challenge due to restricted data access. In this paper, we introduce CADRE (Customizable Assurance of Data REadiness) for FL, a novel framework that allows users to define custom data readiness (DR) standards, metrics, rules, and remedies tailored to specific FL tasks. Our framework generates comprehensive DR reports based on the user-defined metrics, rules, and remedies to ensure datasets are optimally prepared for FL while preserving privacy. We demonstrate the framework’s practical application by integrating it into an existing PPFL framework. We conducted experiments across six diverse datasets, addressing seven different DR issues. The results illustrate the framework’s versatility and effectiveness in ensuring DR across various dimensions, including data quality, privacy, and fairness. This approach enhances the performance and reliability of FL models as well as utilizes valuable resources by identifying and addressing data-related issues before the training phase.
@@
Federated learning, Data readiness for AI, data quality assessment
@@ I Introduction
Federated Learning (FL) [1, 2] is a decentralized machine learning approach allowing multiple participants to train a model collaboratively without sharing their raw data. Rather than centralizing data, FL allows each participant to locally train a model on their data and transmit only the model updates to a central server. This method enhances privacy and security by keeping sensitive data on local devices. However, new challenges emerge when privacy-preserving techniques are applied in FL. A recent study [3] (part of a series on Privacy-Preserving Federated Learning (PPFL) by NIST in collaboration with the UK government’s Responsible Technology Adoption Unit) highlights significant challenges, primarily due to its unique approach that prevents organizations from accessing training data. This restriction complicates essential pre-processing tasks like data cleaning and feature selection, as data scientists cannot view data across different sites. This may lead to potential inconsistencies and deployment failures. Many studies [4, 5, 6] in the literature have demonstrated that low-quality data directly impacts the model by significantly lowering the performance and robustness. Additionally, PPFL’s privacy protections make it difficult to detect poor-quality or maliciously crafted data, which may lead to degrading the final model’s quality. While recent research is beginning to address these issues with techniques like secure input validation and adaptations of data poisoning defenses [7, 8], these solutions are not yet widely implemented in practical PPFL libraries.
@@
In our efforts to address these challenges in PPFL, we introduce Data Readiness for AI (DRAI) into the PPFL domain. Our recent survey [9] presented a comprehensive six-pillar taxonomy for assessing DRAI, focusing on data quality, organization, fairness, understandability, governance, and value. To operationalize these dimensions, we developed AIDRIN (AI Data Readiness Inspector) [10], a framework designed to evaluate the DRAI of datasets across these pillars. However, AIDRIN was initially designed for centralized AI training, where data is uploaded to a standalone platform for evaluation. In contrast, PPFL requires a decentralized approach to data readiness (DR) assessment including methods that preserve privacy and security while evaluating DR across distributed datasets. In addition AIDRIN was designed to visualize the metric evaluations that were predefined. This study aims to bridge the gap in meeting specific DR standards tailored to PPFL tasks by allowing users to define custom metrics and evaluation criteria to suit the FL context while not compromising privacy.
@@
An extendable framework for supporting user-defined metrics, rules, and remedies is still unavailable. For example, in the healthcare industry, PPFL can be used to develop a predictive model for diagnosing a specific disease using MRI scans from multiple hospitals [11]. However, challenges such as data heterogeneity, data quality, and privacy concerns arise. Hospitals often use different MRI machines, which leads to variations in image quality, resolution, and file format due to differences in hardware, software, and imaging protocols. In addition, some datasets may contain noisy or incomplete images, caused not only by machine differences but also by factors such as scanning artifacts, acquisition errors, or data corruption. To address these challenges we can allow stakeholders to define custom DR standards, metrics, rules, and remedies tailored to this FL task. Experts can establish standards for what constitutes an “AI-ready” MRI scan, such as specific format and resolution requirements, and implement metrics to evaluate data quality. Rules can be set to automatically flag images that do not meet these standards, and remedies, such as pre-processing techniques, can be applied to improve data quality. It is required that each hospital to ensure independently that its data meets the necessary standards before participating in the FL process. This may result in a consistent and high-quality dataset across all participants. Consequently, the reliability and performance of the predictive model are improved while maintaining patient privacy.
@@
To meet these challenging requirements in preparing and ensuring DR in PPFL, we propose a novel framework, called CADRE (Customizable Assurance of Data REadiness). This framework allows FL users to define custom DR standards, such as metrics, rules, and remedies tailored to specific FL tasks. In this context, “users” refers to the individuals or stakeholders who are involved in and responsible for a specific FL task, and who collaborate to establish the necessary standards, metrics, rules, and remedies. In the rest of the paper, “users” will be used to denote these key participants. This framework allows clients to locally execute these user-defined functions to dynamically ensure their data meets the necessary standards without compromising privacy. Clients can verify compliance with these rules and apply remedies to their data if necessary. The results of these metric evaluations are compiled into a DR report for user inspection. The report includes evaluations based on the custom readiness standards, along with standard metrics and visualizations of client data statistics. This framework brings a human-in-the-loop approach to FL by involving users in the definition, validation, and refinement of DR standards.
@@
This process ensures that only clients with qualified data participate in the FL system. By doing so, it maintains DR and integrity while preserving privacy. The framework is designed to be generalizable, applicable to any FL task, and adaptable to various domains. To demonstrate its practical application, we have developed an extensible module for the APPFL (Advanced Privacy-Preserving Federated Learning) framework [12, 13], an open-source software framework that enables researchers and developers to implement, test, and validate various PPFL techniques. By this integration, we showcase how this approach can be used in existing PPFL workflows. The main contributions of this study are highlighted below.
  1. 1.
@@
    We propose a novel framework that enables users within a PPFL system to define custom metrics, rules, and remedies. Our framework addresses the challenge of executing these user-defined standards by automating the process and ensuring that clients can locally apply these actions to meet required data standards while preserving privacy.
  2. 2.
@@
    The framework generates comprehensive DR reports that evaluate the standards defined by users. This method ensures privacy is preserved by only including aggregated metric evaluations without exposing any raw data. Users can review these reports to assess whether clients have met the expected standards and gain insights into the data’s characteristics.
  3. 3.
@@
    Integrating CADRE into the APPFL framework demonstrates its practical value and compatibility with existing PPFL workflows.
@@
We evaluated CADRE using six datasets with various data modalities (e.g. 2D images, tabular data, 3D volumetric data) and downstream tasks (such as classification, segmentation, and survival analysis). These datasets were either naturally occurring or intentionally constructed to reflect seven key DR challenges: noise, class imbalance, duplicate records, high memory consumption, bias, outliers, and insufficient anonymity. As our framework allows users within a PPFL system to define custom metrics, rules, and remedies, these issues were effectively addressed. Through these experiments, we demonstrated how datasets that initially failed to meet these standards could be systematically improved. This illustrates the versatility and effectiveness of our framework in real-world and heterogeneous data scenarios.
@@ II Related Work
Users seeking to adopt AI rely on structured frameworks to evaluate their DR, with a focus on key aspects such as data quality, governance, and infrastructure. Existing frameworks [14, 15, 16, 17] primarily assess data availability, volume, quality, governance, and ethics. A wide range of data cleansing tools [17, 18, 19] are available today, each offering unique features to ensure the accuracy, reliability, and trustworthiness of data. These tools focus on user-friendly interfaces, advanced profiling, duplicate removal, and other validation rules, enabling both technical and non-technical users to efficiently clean and standardize data.
Despite their strengths, these frameworks exhibit critical gaps when applied to modern, distributed AI environments. Most notably, they lack integration with FL architectures, which are increasingly important as organizations move toward decentralized data models. Existing frameworks generally assume a centralized data environment and provide little to no guidance on assessing DR across distributed clients with heterogeneous data distributions. They also fall short in addressing compliance challenges related to cross-border data flows, which are common in FL scenarios. FL has become a prominent decentralized machine learning approach, allowing multiple clients to collaboratively train models while maintaining data privacy. Apart from its advantages, FL faces key challenges such as robustness against malicious clients and heterogeneous data distributions [12]. Several frameworks have been developed to address these challenges with unique methodologies and trade-offs.
@@
The performance of FL models is often affected due to heterogeneous and noisy data distributions. In FL, where data is distributed between multiple clients, label noise refers to incorrect or inconsistent labels in the training data, which can significantly reduce model performance. To address this issue, FedELC [20] proposes a two stage framework that first identifies clients with noisy labels and then applies label correction strategies to improve the overall robustness of the global model. However, it focuses solely on label noise and ignores other critical aspects of DR. FedDQA [21] introduces a metric to evaluate client data quality without additional computational cost, allowing the selection of higher-quality clients for training. Although effective in minimizing the influence of noisy data, this approach risks introducing selection bias and does not actively improve the underlying data. In the domain of PPFL, methods such as lazy influence approximation [22] and FedDQC [23] offer quality assessments that preserve privacy using influence scores and relevance alignment, respectively. Although these approaches maintain confidentiality, they have computational overhead and suffer from reduced data resolution under strict privacy constraints.
Another key limitation of existing FL frameworks is their lack of flexibility in supporting custom DR metrics and remediation workflows. Most rely on static, predefined evaluation criteria, making it difficult to accommodate domain-specific requirements. Remediation processes are often rigid and lack support for user-specific operations such as federated anonymization or edge-device preprocessing. Even unified data platforms rarely allow integration of custom rules or remedies. To address these gaps, our proposed framework enables users to define customized metrics, rules, and remedies aligned with the needs of specific FL systems. This flexibility helps manage data heterogeneity by enforcing consistent standards across clients, all while preserving privacy. The framework integrates seamlessly with existing PPFL workflows and supports DR evaluation before initiating resource-intensive training. Moreover, it aligns with the vision of Industry 5.0 [24], emphasizing human-centric, privacy-aware, and adaptable AI systems that empower users to take control of DR.
@@ III Design Overview
The objective of CADRE is to allow users of FL systems to define and utilize both foundational and customizable actions. To support this requirement, CADRE provides the following main components: metrics, DR reports, rules, and remedies. The functionality of each of these components will be discussed in details in the following subsections. In Figure 1, we show an outline of CADRE by illustrating its components.
Figure 1: An overview of CADRE framework for FL tasks. Metrics include commonly known standard DR evaluation measurements. The extensible functions are used to define custom DR metrics, rules, and remedies. The DR report consolidates the standard and custom metric evaluations with visualizations.
@@ III-A Metrics Component
We divided the metrics component of CADRE into two main parts: standard metrics and customizable metrics. The standard metrics include a set of universally applicable measures such as evaluating sample sizes, data sparsity, and basic statistical measures like mean, median, and standard deviation of the client’s data distribution. These metrics serve as a baseline for assessing DR of clients’ data across any FL task. Additionally, the standard metrics module contains basic visualizations, such as bar charts and scatter plots, which are included in the DR reports to provide a visual representation of the client data’s characteristics.
The customizable metrics sub-module offers an extensible function that allows users to define custom metrics tailored to their unique FL task and evaluation needs. This flexibility ensures that users can assess client data according to the specific requirements of their projects. For example, if a particular task requires assessing the completeness or skewness of the data, users can define these metrics within CADRE. These standard and custom metric evaluations and visualizations allow users to quickly grasp the readiness of clients’ data and identify potential issues that may lead to unexpected behavior in downstream FL tasks.
III-B Rules and Remedies Sub-Modules
Beyond metrics, CADRE includes a robust sub-module for defining rules and remedies. This sub-module allows users to establish custom rules that data must meet to be considered ready for the next stages of the FL pipeline. The users can also define custom remedies to improve the readiness of the data to meet the specified rules.
For instance, if users need to assess noise levels in the data, they can use metrics such as the standard deviation of the data distribution to quantify noise. A high standard deviation may indicate excessive variability and suggest the presence of noise. The users can then establish a rule where the standard deviation must not exceed a predefined threshold. If this threshold is surpassed, remedies could be implemented, such as filtering out extreme values or including only a subset of the affected client’s data in the analysis.
@@ III-C DR Reporting Module
CADRE generates detailed DR reports by aggregating metric evaluations and visualizations produced by individual clients. It also includes principal component analysis (PCA) [25] graphs, to illustrate the combined data distribution and highlight the heterogeneity among clients. These insights are compiled into an easily readable HTML report, allowing users to quickly assess whether clients meet specified standards while ensuring data privacy. This feature is essential to maintain transparency and accountability throughout the DR process.
For instance, for a given FL task, a custom metric could involve measuring class imbalance within each client’s dataset in the FL system. Identifying class imbalance is important because it can bias the learning process, especially in classification tasks where underrepresented classes may be poorly learned [26]. In this scenario, the rule would be to flag any client datasets where the class distribution significantly deviates from a defined threshold of balance. If a client is flagged, the remedy might involve data augmentation or re-sampling techniques to mitigate the imbalance until the metric indicates an acceptable distribution. The resulting report will display the class distribution statistics for each client’s dataset, making it easy to identify and address any flagged issues. Figure 2 presents an actual HTML DR report generated for this specific example during an FL experiment. The report includes evaluations of both standard and custom metrics, visualizations for each of the two clients involved in the experiment, and combined plots. The visualizations include standard plots such as class distribution and data distribution charts, while the combined plot is a PCA visualization of a sample of the data from the clients. For this example, we used the Adult Income dataset [27], and CADRE is integrated into the APPFL framework. More details about this integration and the experiments can be found in sections IV and V.
@@
Clients participating in the FL framework use custom metrics within CADRE to locally evaluate their data and generate DR reports. If the client data meets the specified rules, the data will proceed to the subsequent stages of the FL pipeline. Conversely, if the data does not meet the rules, remedies defined by the users within CADRE will be applied to improve the DR. This process will iterate until the data complies with the established rules. This will ensure DR for the next stages of the FL pipeline. Figure 3 provides a visual representation of this iterative approach by illustrating how clients use CADRE’s functions to assess DR, apply custom rules, and implement remedies as needed while preserving privacy.
Figure 3: The figure illustrates the iterative data evaluation and remediation process within the clients involved in FL framework. It outlines how clients use the CADRE’s functions to assess DR, apply custom rules, and implement remedies as needed.
By integrating these sub-modules, CADRE provides a comprehensive and flexible framework for ensuring DR in FL systems. This framework allows users to tailor the DR process to the specific needs of their projects while maintaining high standards of DR and privacy.

### v2

@@ Abstract
Privacy-Preserving Federated Learning (PPFL) is a decentralized machine learning approach where multiple clients train a model collaboratively. PPFL preserves the privacy and security of a client’s data without exchanging it. However, ensuring that data at each client is of high quality and ready for federated learning (FL) is a challenge due to restricted data access. In this paper, we introduce CADRE (Customizable Assurance of Data REadiness) for federated learning (FL), a novel framework that allows users to define custom data readiness (DR) metrics, rules, and remedies tailored to specific FL tasks. CADRE generates comprehensive DR reports based on the user-defined metrics, rules, and remedies to ensure datasets are prepared for FL while preserving privacy. We demonstrate a practical application of CADRE by integrating it into an existing PPFL framework. We conducted experiments across six datasets and addressed seven different DR issues. The results illustrate the versatility and effectiveness of CADRE in ensuring DR across various dimensions, including data quality, privacy, and fairness. This approach enhances the performance and reliability of FL models as well as utilizes valuable resources.
@@
Data readiness for AI, Data quality assessment, Federated learning,
@@ I Introduction
Federated Learning (FL) [1, 2] allows multiple decentralized participants to train a model collaboratively without sharing their raw data. Rather than centralizing data, FL allows each participant to locally train a model on their data and transmit only the model updates to a central server. This method enhances privacy and security by keeping sensitive data locally. However, new challenges emerge when privacy-preserving techniques are applied in FL. A recent study on Privacy-Preserving Federated Learning (PPFL) led by NIST [3] highlights significant challenges, primarily due to the lack of access to training data. Data cleaning and feature selection are complicated because data scientists cannot view data across different clients. This may lead to inconsistencies and deployment failures. Many studies [4, 5, 6] have demonstrated that low-quality data directly impacts the model by lowering the performance and robustness. Additionally, PPFL’s privacy protections make it difficult to detect poor-quality or maliciously crafted data, which may lead to degrading the final model’s quality. While recent research is beginning to address these issues with techniques like secure input validation and adaptations of data poisoning defenses [7, 8], these solutions are not yet widely implemented in practical PPFL libraries.
@@
In our efforts to address these challenges, we introduce Data Readiness for AI (DRAI) into the PPFL domain. Our recent survey [9] presented a comprehensive taxonomy for assessing DRAI, focusing on data quality, organization, fairness, understandability, governance, and value. We developed AIDRIN (AI Data Readiness Inspector) [10] framework to evaluate the DRAI of datasets across these dimensions. However, AIDRIN was initially designed for centralized AI training, where data is uploaded to a standalone platform for evaluation. In contrast, PPFL requires decentralized data readiness (DR) assessment, including methods that preserve privacy and security.
@@
A framework for supporting user-defined metrics, rules, and remedies to allow data stewards and FL administrators to define custom metrics and evaluation criteria while preserving privacy is needed. However, such a framework targeting either PPFL or centralized data readiness assessment is still unavailable. For example, in healthcare, PPFL can be used to develop a model for diagnosing a specific disease using MRI scans from multiple hospitals [11]. However, challenges such as data heterogeneity, quality, and privacy concerns arise because hospitals often use different MRI machines that leads to variations in image quality, resolution, and file format due to differences in hardware, software, and imaging protocols. In addition, some datasets may contain noisy or incomplete images, caused not only by machine differences but also by factors such as scanning artifacts, acquisition errors, or data corruption. To address these challenges, data owners should define custom DR standards, metrics, rules, and remedies tailored to their FL tasks. Data owners can establish standards for what constitutes an “AI-ready” MRI scan, such as data format and resolution requirements, and implement metrics to evaluate quality. Rules can be set to automatically flag images that do not meet these standards, and remedies, such as pre-processing techniques, can be applied to improve quality. It is required that each hospital ensure independently that its data meets the standards before participating in the FL process.
@@
To meet these challenging requirements in preparing and ensuring DR in PPFL, we propose a novel framework, called CADRE (Customizable Assurance of Data REadiness). This framework allows FL “administrators” to define custom data readiness (DR) standards, including metrics, rules, and remedies tailored to specific FL tasks. Here, administrators refers to the individuals or stakeholders responsible for a given FL task who collaborate to establish these data readiness standards. CADRE allows clients to locally execute these custom functions to ensure their data meets the necessary standards at run time without compromising privacy. Clients can verify compliance with these rules and apply remedies to their data as necessary. The results of these metric evaluations are compiled into a DR report for administrators’ inspection. The report includes evaluations based on the custom readiness standards, along with standard metrics and visualizations of client data statistics. This framework brings a human-in-the-loop approach to FL by involving administrators in the definition, validation, and refinement of DR standards. CADRE can include predefined techniques and rules to showcase its capabilities, but its primary functionality lies in its customizability. The framework enables administrators to define a wide range of DRAI evaluation metrics, rules, and remedies tailored to specific FL tasks. This flexibility makes CADRE adaptable and practical for diverse PPFL scenarios, which enhances its usefulness across applications.
@@
DR evaluation ensures that only clients with qualified data participate in the FL system. The CADRE framework is designed to be generalizable, applicable to any FL task, and adaptable to various domains. To demonstrate its practical application, we have developed an extensible module for the APPFL (Advanced Privacy-Preserving Federated Learning) framework [12, 13], an open-source software framework that enables researchers and developers to implement, test, and validate various PPFL techniques. With this integration, we showcase usage of CADRE in existing PPFL workflows. The main contributions of this study are:
@@
  • We propose a novel framework that enables FL administrators within a PPFL system to define custom metrics, rules, and remedies. CADRE addresses the execution of these custom standards by automating the process and ensuring that clients can locally apply these actions to meet required data standards while preserving privacy.
@@
  • We generate comprehensive DR reports in CADRE that evaluate the metrics defined by FL administrators. This ensures privacy preservation by only including aggregated metric evaluations without exposing any raw data. Administrators can review these reports to assess whether clients have met the expected standards and gain insights into the data’s characteristics.
@@
  • We integrate CADRE into APPFL, demonstrating compatibility with existing PPFL workflows.
@@
We evaluated CADRE using six datasets with a variety of data modalities (e.g., 2D images, tabular data, 3D volumetric data) and downstream tasks (such as classification, segmentation, and survival analysis). In some cases, we polluted the datasets to add noise, class imbalance, duplicate records, high memory consumption, bias, outliers, and insufficient anonymity. CADRE allows administrators within a PPFL system to define custom metrics, rules, and remedies, showing that the issues caused by our pollution were effectively addressed. We also demonstrate CADRE’s impact further using an example where resolving DR challenges leads to improvements in model performance. In the remainder of the paper, we describe related work (§II), CADRE design (§III), its integration into APPFL (§IV), and its evaluation (§V).
@@ II Related Work
A few frameworks evaluate data with a focus on aspects such as data quality, governance, and infrastructure. Existing frameworks [14, 15, 16, 17] primarily assess data availability, volume, quality, governance, and ethics. A wide range of data cleansing tools [17, 18, 19] are available today, each offering unique features to ensure the accuracy, reliability, and trustworthiness of data. However, most users prefer manual cleaning of the data and decide on AI readiness themselves or skip these tools entirely.
Despite their strengths, these frameworks exhibit critical gaps when applied to modern, distributed AI environments. They lack integration with FL architectures. Existing frameworks generally assume a centralized data environment. They also fall short in addressing compliance challenges related to cross-border data flows, which are common in FL scenarios.
@@
The performance of FL models is often affected due to heterogeneous and noisy data distributions. In FL, where data is distributed between multiple clients, label noise refers to incorrect or inconsistent labels in the training data, which can significantly reduce model performance. To address this issue, FedELC [20] proposes to identify clients with noisy labels and apply label correction strategies to refine the labels. However, it ignores other critical aspects of DR. FedDQA [21] introduces a metric to evaluate client data quality, allowing the selection of higher-quality clients for training. Although effective in minimizing the influence of noisy data, this approach risks introducing selection bias and does not actively improve the underlying data. In the domain of PPFL, methods such as lazy influence approximation [22] and FedDQC [23] offer quality assessments that preserve privacy using influence scores and relevance alignment, respectively. Although these approaches maintain confidentiality, they have computational overhead and suffer from reduced data resolution under strict privacy constraints.
Another key limitation of existing FL frameworks is their lack of flexibility in supporting custom DR metrics and remediation workflows. Most rely on static, predefined evaluation criteria, making it difficult to accommodate domain-specific requirements. Remediation processes are often rigid and lack support for custom operations such as federated anonymization or edge-device preprocessing. Even unified data platforms rarely allow integration of custom rules or remedies. To address these gaps, our proposed framework enables FL administrators to define customized metrics, rules, and remedies aligned with the needs of specific FL systems. This flexibility helps manage data heterogeneity by enforcing consistent standards across clients, all while preserving privacy. The framework integrates seamlessly with existing PPFL workflows and supports DR evaluation before initiating resource-intensive training. Moreover, it aligns with the vision of Industry 5.0 [24], emphasizing human-centric, privacy-aware, and adaptable AI systems that empower administrators to take control of DR.
@@ III Design Overview
The objective of CADRE is to allow administrators of FL systems to define and utilize both foundational and customizable actions. To support this, CADRE provides the following main components: metrics, DR reports, rules, and remedies. In Figure 1, we show an outline of CADRE with its components, including metrics that are standard and custom (i.e., administrator-defined). Rules and remedies are also administrator-defined functions.
Figure 1: An overview of CADRE framework for FL tasks. Metrics include commonly known standard DR evaluation measurements. The extensible functions are used to define custom DR metrics, rules, and remedies. The DR report provides standard and custom metric evaluations with visualizations.
@@ III-A Metrics Component
We divided the metrics component of CADRE into two main parts: standard metrics and custom metrics. The standard metrics include a set of DR metrics defined in AIDRIN [10] including quality metrics such as evaluating sample sizes, data sparsity, and statistical measures like mean, median, and standard deviation of the client’s data distribution. These metrics serve as a baseline for assessing DR of clients’ data across any FL task. Additionally, the standard metrics component contains basic visualizations, such as bar charts and scatter plots, which are included in the DR reports to provide a visual representation of the client data characteristics.
The custom metrics component is an extensible capability that allows FL administrators to provide custom metrics tailored to their unique FL task and evaluation needs. This flexibility ensures that administrators can assess client data according to the specific requirements of their projects. For example, if a task requires assessing the completeness or skewness of the data, administrators can define these metrics within CADRE. These standard and custom metric evaluations and visualizations allow administrators to quickly grasp the readiness of clients’ data and identify potential issues that may lead to unexpected behavior in downstream FL tasks.
III-B Rules and Remedies Components
Besides metrics, CADRE also includes sub-modules for defining rules and remedies. These sub-modules allow administrators to establish custom rules that their custom metric must meet to be considered ready for the next stages of the FL pipeline. The administrators can also define custom remedies to improve the readiness of the data to meet the specified rules. For instance, if administrators need to assess noise levels in the data, they can use metrics such as the standard deviation of the data distribution to quantify noise. A high standard deviation may indicate excessive variability and suggest the presence of noise. The administrators can then establish a rule where the standard deviation must not exceed a predefined threshold. If this threshold is surpassed, remedies could be implemented, such as filtering out extreme values or including only a subset of the affected client’s data in the analysis.
@@ III-C DR Reporting Module
CADRE generates detailed DR reports by aggregating metric evaluations and visualizations produced by individual clients. It also includes principal component analysis (PCA) [25] graphs, to illustrate the combined data distribution and heterogeneity among clients. These insights are compiled into an easily readable HTML report, allowing administrators to assess whether clients meet specified standards while ensuring data privacy. This feature is essential to maintain transparency and accountability throughout the DR process.
For instance, for a given FL task, a custom metric could involve measuring class imbalance within each client’s dataset in the FL system. Identifying class imbalance is important because it can bias the learning process, especially in classification tasks where underrepresented classes may be poorly learned [26]. In this scenario, a rule would be to flag any client datasets where the class distribution significantly deviates from a defined threshold of balance. If a client is flagged, the remedy might involve data augmentation or re-sampling techniques to mitigate the imbalance until the metric indicates an acceptable distribution. The resulting report will display the class distribution statistics for each client’s dataset, making it easy to identify and address any flagged issues. In Figure 2, we present a DR report generated for this specific example. The report includes evaluations of both standard and custom metrics, visualizations for each of the two clients involved in the experiment, and combined plots. The visualizations include standard plots such as class distribution and data distribution charts, while the combined plot is a PCA visualization of a sample of the data from the clients. For this example, we used the Adult Income dataset [27], and CADRE is integrated into the APPFL framework. More details about this integration and the experiments can be found in sections IV and V.
@@
Clients participating in the FL framework use custom metrics within CADRE to locally evaluate their data and generate DR reports. If the client data meets the specified rules, the data will proceed to the subsequent stages of the FL pipeline. Conversely, if the data does not meet the rules, remedies defined by the administrators within CADRE will be applied to improve the DR. This process will iterate until the data complies with the established rules. This will ensure DR for the next stages of the FL pipeline. Figure 3 provides a visual representation of this iterative approach by illustrating how clients use CADRE’s functions to assess DR, apply custom rules, and implement remedies while preserving privacy.
Figure 3: An iterative data evaluation and remediation process where clients are involved in an FL framework. It outlines how clients use the CADRE’s functions to assess DR, apply custom rules, and implement remedies as needed.
By integrating metrics, rules, and remedies components into DR frameworks, CADRE provides a comprehensive and flexible framework for ensuring DR in FL systems. This framework allows administrators to tailor the DR process to the specific needs of their projects while maintaining high standards of DR and privacy.